OpenStudy (anonymous):

Is 38/53 in simplest form? Explain why or why not. If not write it in simplest form. PLEASE HELP WILL MEDAL!

OpenStudy (anonymous):

@cathyangs @Cherryboa

OpenStudy (anonymous):

@hhmorris98 @stuck-help

OpenStudy (anonymous):

@MW3030

OpenStudy (anonymous):

@leahhhmorgannn

OpenStudy (anonymous):

@superdavesuper

OpenStudy (cathyangs):

Yes. 53 is prime so there is no simplification.

OpenStudy (anonymous):

I believe that its in its most simplest form

OpenStudy (anonymous):

you cant simplify it anymore than it is

OpenStudy (anonymous):

I need 3 good sentences to explain why.

OpenStudy (anonymous):

@campbell_st @arakacombs @themazinmathmaster

OpenStudy (anonymous):

go for prime factors top and bottom 38 = 2 * 19 53 = 53 * 1 nothing to cancel in 2 * 19/53 so 38/53 IS as simple as you can write it.

OpenStudy (campbell_st):

start with the fact that 53 is a prime number. And only has factors of...

OpenStudy (anonymous):

none

OpenStudy (anonymous):

1

OpenStudy (anonymous):

53

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Can you help somebody?

OpenStudy (anonymous):

Thank you everyone!

OpenStudy (cathyangs):

So because 53 is a prime, its only factors are 1 and 53. 38 can be factored to 2 * 19. They share no common factors, so it cannot be simplified further; thus 38/53 is the simplest form.

OpenStudy (anonymous):

do you still need help?